OpenStudy (anonymous):

I'm trying to graph an equation by plotting the points but first I must find the missing point. Can anyone show me how to do this? The problem is y=1/5x-3 and rhe points are (5,_) and (0,_).

OpenStudy (anonymous):

Plug the x values (5 and 0) into the equation given to get the y values for each of those points.

OpenStudy (anonymous):

points are in form of (x,y) so your point there, you can substitute the first number as x into the equation, so you can find y. Do the same for the second point

OpenStudy (amistre64):

find your y and x intercepts and draw the line :)

OpenStudy (amistre64):

y = (1/5)(0) - 3 what does y =

OpenStudy (amistre64):

0 = (1/5)x -3 what will x = or if you want to find that arrbitrary point of (5,y) y = 5/5 - 3 y =

OpenStudy (anonymous):

thankx